A prospective MBA student earns $60,000 per year in her current job and expects that amount to increase by 9% per year. She is considering leaving her job to attend business school for two years at a cost of $50,000 per year. She has been told that her starting salary after business school is likely to be $85,000 and that amount will increase by 13% per year. Consider a time horizon of 10 years, use a discount rate of 15%, and ignore all considerations not explicitly mentioned here.

Assume all cash flows occur at the start of each year (i.e., immediate, one year from now, two years from now,..., nine years from now). Also assume that the choice can be implemented immediately so that for the MBA alternative the current year is the first year of business school.

What is the net present value of the more attractive choice?

Please round your answer to the nearest dollar.
Answer #1

Current Salary Per year = 60,000
Growth =9%
Time = 10 years
PV of Current Job Salary = 60,000 + 60000*(1+9%)/(1+15%) + 60000*(1+9%)2/(1+15%)2+ ........60000*(1+9%)9/(1+15%)9
= 60000* ((1.09/1.15)10-1)/(1.09/1.15-1) = 477047.49
PV of the MBA fees = 50000+50000/(1+15%)1 = 93,478.26

PV of Salary after MBA = 65000+ 65000*(1+13%)/(1+15%)+......65000*(1+13%)/(1+15%)7
= 60000*((1.13/1.15)8-1)/(1.13/1.15-1) = 489425.17

PV of Salary at present = 489425.17/1.152 = 370,075.74

NPV = 370,075.74 - 477,047.49 - 93,478.26 = -200450.01

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well